We have detected a fiber break on a span between TH2 and DC5.
Span between DC5 and DC3 is still up and traffic still flows through this optical path with 400G capacity.
We are investigating with our dark fiber provider.
This concerns both public and private (RPN) network.
05/06/19 1510Z (1710LT)
Provider's maintenance still ongoing.
We don't have any specific ETA yet, we will keep monitoring the situation.